Crisp Cookware's common stock is expected to pay a dividend of $1.75 a share at the end of this year (D1 = $1.75); its beta is 0.70; the risk-free rate is 3.5%; and the market risk premium is 4%. The dividend is expected to grow at some constant rate g, and the stock currently sells for $42 a share. Assuming the market is in equilibrium, what does the market believe will be the stock's price at the end of 3 years (i.e., what is  )? Do not round intermediate steps. Round your answer to the nearest cent.

$  
Answer #1

Expected return = risk free rate + beta * market risk premium

= 3.5% + 0.7 * 4%

= 6.3%

Price = dividend next year/(Required return - growth rate)

=>

42 = 1.75/(0.063 - g)

=>

growth rate g = 2.13%

Price in year 3 = D4/(r-g)

= 1.75 * 1.0213^3/(0.063-0.0213)

= 44.71